OpenStudy (anonymous):

I REALLY NEED HELP ON THIS ONE QUESTION!!! WILL FAN AND MEDAL IF HELPED!! I am supposed to graph some points to see where they intersect but I don't know which site to do it on or how exactly because I tried and it didn't work.

OpenStudy (anonymous):

This is the question: Graph the system of equations and show the point of intersection. 4 192 5 186.62 4 215 5 223.95 These are the points, both on week 4 through 5

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@jennyrlz

OpenStudy (anonymous):

@Butterfield1215

OpenStudy (butterfield1215):

idk sorry

OpenStudy (anonymous):

That's okay, lol :/

OpenStudy (anonymous):

I will try to help. Ok

OpenStudy (anonymous):

Okay haha thank you!

OpenStudy (anonymous):

Are those the points?

OpenStudy (anonymous):

I just don't know how to graph the points so they are intersecting

OpenStudy (anonymous):

Yea, I have the points for weeks 4 and 5 for two different equations

OpenStudy (anonymous):

In a biological lab, the cell growth rate of two different organisms is tracked and recorded each week. Given the growth rate, the number of organisms can be determined using the following equations: s(x) = 100 + 23x m(x) = 90(1.2x)

OpenStudy (anonymous):

Ok, give me a minute.

OpenStudy (anonymous):

that's the full question, and I had to find out the values for weeks 4 and 5 which I did but now I have to graph them...

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So you want me to graph this 4 192 5 186.62 4 215 5 223.95

OpenStudy (anonymous):

Yes that's what I'm supposed to graph, and the points are supposed to intersect

OpenStudy (anonymous):

Do you want the 4's to intersect the 5's?

OpenStudy (anonymous):

I think so, the question didn't really say how I was supposed to graph them

OpenStudy (anonymous):

The intersection point is (4.83, 204.18)

OpenStudy (anonymous):

I think that if they intersect that's good enough because I tried to graph them in different ways and they wouldn't cross each other

OpenStudy (anonymous):

Is there a graph to go with that?

OpenStudy (anonymous):

I already gave you the intersection points.

OpenStudy (anonymous):

Sure, give me a minute and I will upload a screenshot

OpenStudy (anonymous):

Yea but my problem asks for me to give a graph so they can see the intersection points

OpenStudy (anonymous):

okay thank you!

OpenStudy (anonymous):

Here

OpenStudy (anonymous):

I graphed it, and labeled the intersection point "E"

OpenStudy (anonymous):

Okay thank you a lot!!! That helped a lot :)

OpenStudy (anonymous):

I will give you a fan and medal if you do the same for me

OpenStudy (anonymous):

Already did, haha :)

OpenStudy (anonymous):

I gave youa fan and medal! Bye

OpenStudy (anonymous):

haha okay thanks!! bye :)